\begin{thebibliography}{99}


\bibitem {Baker} J. A. Baker, 
\newblock{\it The stability of certain functional equations},
 \newblock {Proc. Amer. Math. Soc.} {\bf 112} (1991), 729--732. 

 
\bibitem {Cadariu1} L. Cadariu, V. Radu, 
\newblock{\it Fixed points and the stability of Jensesn's functional equation},
 \newblock {J. Inequal. Pure Appl. Math.} {\bf 4} (2003), no. 1, Art. No. 4, 1--7.

\bibitem {Cadariu2} L. Cadariu, V. Radu,  
\newblock{\it Fixed point methods for the generalized stability of functional equations in a single variable},
 \newblock {Fixed Point Theory Appl.} {\bf 2008} (2008), Art. ID 749392.



\bibitem {Ali} A. Ebadian, I. Nikoufar, Th. M. Rassias, G. Ghobadipour, 
\newblock{\it Stability of generalized derivation on Hilbert $C^{*}$-modules associated with a Pexiderized Cauchy-Jensen type functional equation},
 \newblock {Acta Math. Sci.} {\bf 32} (2012), 1226--1238.
%
%\bibitem {eshaghi1} M. E. Gordji, A. Najati, A. Ebadian,
%\newblock{\it Stability and superstability of Jordan homomorphisms and Jordan derivations on Banach algebras and $C^{*}$-algebras: A fixed point approach},
% \newblock {Acta Math. Sci.} {\bf 31B} (2011), 1911--1922.
%%%
%\bibitem {eshaghi2} M. E. Gordji, C. Park, M. B. Savadkouhi, 
%\newblock{\it The stability of quartic functional equation with the fixed point alternative},
% \newblock {Fixed Point Theory.} {\bf 11} (2010), 265--272.


\bibitem {Gajda} Z. Gajda,
\newblock{\it On stability of additive mappings},
 \newblock {Int. J. Math. Math. Sci.} {\bf 14} (1991), no. 1, 431--434.

\bibitem {Gavruta} P. G\v{a}vruta, 
\newblock{\it A generalization of the Hyers-Ulam-Rassias stability of approximately additive mappings},
 \newblock {J. Math. Anal. Appl.} {\bf 184} (1994), 431--436.

\bibitem {Hyers} D. H. Hyers,
\newblock{\it On the stability of the linear functional equation},
 \newblock {Proc. Natl. Acad. Sci. U.S.A.} {\bf 27} (1941), 222--224.

\bibitem {Isac} G. Isac, Th. M. Rassias,
\newblock{\it Stability of $\Psi$-additive mappings: Applications to nonlinear analysis},
 \newblock {Int. J. Math. Math. Sci.} {\bf 19} (1996), 219--228.


\bibitem {kaplansky} I. Kaplansky,
\newblock{\it Modules over operator algebras},
 \newblock {Amer. J. Math.} {\bf 75} (1953), no. 4, 839--858.



\bibitem {Sajjad1} S. Khan, S. W. Min, C. Park,  
\newblock{\it Permuting triderivations and permuting trihomomorphisms in complex Banach algebras},
 \newblock {J. Anal.} {\bf 32} (2024), 1671--1690.


\bibitem {Sajjad3} S. Khan, S. W. Min, C. Park,  
\newblock{\it Hyers-Ulam stability of fuzzy Hilbert $C^{*}$-module homomorphisms and fuzzy Hilbert $C^{*}$-module derivations}, preprint.


\bibitem {Sajjad4} S. Khan, S. W. Min, C. Park,  
\newblock{\it Stability of Jordan $\ast$-homomorphisms in Hilbert $C^{*}$-modules}, preprint.




\bibitem {Sajjad2} S. Khan, C. Park, M. Donganont,
\newblock{\it The stability of bi-derivations and bihomomorphisms in Banach algebras},
 \newblock {J. Math. Comput. Sci.} {\bf 35} (2024), no. 4, 482--491.
%
%\bibitem {Hilbert CSMD} E. C. Lance, 
%\newblock{\it Hilbert $C^*$-Modules: A Tool for Operator Algebras},
%Cambridge University Press, Cambridge, 1995.

\bibitem {Lance} E. C. Lance,
\newblock{\it Hilbert $C^{*}$-Modules: A Tool for Operator Algebras}, Cambridge University Press, Cambridge, 1995.

\bibitem {m-r} D. Mihet, V. Radu, 
\newblock{\it On the stability of the additive Cauchy functional equation in random normed spaces},
 \newblock {J. Math. Anal. Appl.} {\bf 343} (2008), 567--572.



%
%\bibitem {moh} M. Mohammadi, Y. Cho, C. Park, P. Vetro, R. Saadati, 
%\newblock{\it Random stability of an additive-quadratic-quartic functional equation}, 
%\newblock {J. Inequal. Appl.} {\bf 2010} (2010), Article ID 754210, 18 pages.
%
%\bibitem {naj} A. Najati, Y. Cho,   
%\newblock{\it Generalized Hyers-Ulam stability of the Pexiderized Cauchy functional equation  in non-Archimedean spaces},
% \newblock {Fixed Point Theory Appl.} {\bf 2011} (2011), Article ID 309026, 11 pages.

\bibitem {Ismail} I. Nikoufar, 
\newblock{\it Jordan $(\theta , \phi)$-derivations on Hilbert $C^*$-modules},
 \newblock {Indag.  Math.} {\bf 26} (2015), 421--430.

\bibitem{Siriluk} S. Paokanata, M. Dehghanian, C. Park, Y. Sayyari, 
\newblock {\it A system of additive functional equations in complex Banach algebras},
\newblock {Demonstr.  Math.} {\bf 56} (2023),  no. 1, Art. ID 20220165.

\bibitem{c-linear} C. Park,
\newblock{\it Homomorphism between Poisson $JC^*$-algebras},  
\newblock {Bull. Braz. Math. Soc.} {\bf 36} (2005), 79--97.

\bibitem {Park rho 1} C. Park,
\newblock{\it Additive $\rho$-functional inequalities and equations},
 \newblock {J. Math. Inequal.} {\bf 9} (2015), 17--26.


\bibitem {Park rho 2} C. Park,  
\newblock{\it Additive $\rho$-functional inequalities in non-Archimedean normed spaces},
 \newblock {J. Math. Inequal.} {\bf 9} (2015), 397--407.

\bibitem{Park 2017} C. Park,
\newblock{\it Fixed point method for set-valued functional equations},  
\newblock { J. Fixed Point Theory Appl.} {\bf 19} (2017), 2297--2308.

\bibitem{Park bi-derivation} C. Park,
\newblock{\it Biderivations and bihomomorphisms in Banach algebras},  
\newblock {Filomat} {\bf 33} (2019), 2317--2328.



\bibitem{Radu} V. Radu,
\newblock{\it The fixed point alternative and the stability of functional equations},  
\newblock {Fixed Point Theory} {\bf 4} (2003), 91--96.

\bibitem{Rassias 1978} Th. M. Rassias,
\newblock{\it On the stability of the linear mapping in Banach spaces},  
\newblock {Proc. Amer. Math. Soc.} {\bf 72} (1978), 297--300.

\bibitem{Rassias 1990} Th. M. Rassias, 
\newblock{\it Problem 16; 2; Report of the 27th International Symposium on Functional Equations},
\newblock {Aequationes Math.} {\bf 39} (1990), no. 4, 292--293. 
%
%\bibitem{Rassias 1998} Th. M. Rassias, 
%\newblock{\it On the stability of the quadratic functional equation and it's application},
%\newblock {Studia Univ. Babes Bolyai.} {\bf XLIII} (1998), 89--124.

\bibitem{Rassias 1993} Th. M. Rassias, P. \v{S}emrl, 
\newblock{\it On the Hyers-Ulam stability of linear mappings},
\newblock{J. Math. Anal. Appl.} {\bf 173} (1993), 325--338.

\bibitem{Homomorphism definition} H. Saidi, A. R. Janfada, M. Mirzawaziri,
\newblock{\it Kinds of derivations on Hilbert $C^*$-modules and their operator algebras},
\newblock {Miskolc Math. Notes} {\bf 16} (2015), no. 1, 453--461.


\bibitem{Ulam} S. M. Ulam, 
\newblock {\it Problems in Modern Mathematics},
\newblock{John Wiley and Sons, New York,}  1964.


\end{thebibliography}

\documentclass[preprint,moreauthors,11pt]{elsarticle}

\usepackage{amssymb}

\usepackage{lineno,color}
\usepackage{amsmath, amsthm}

\theoremstyle{definition}
  \newtheorem{defn}{Definition}[section]
  
\theoremstyle{plain}
  \newtheorem{lem}[defn]{Lemma}
  \newtheorem{thm}[defn]{Theorem}
  
  \newtheorem{cor}[defn]{Corollary}

\numberwithin{equation}{section}

\usepackage{lineno,hyperref}

\journal{j}

\begin{document}
\begin{frontmatter}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{example}[theorem]{Example}
\newtheorem{xca}[theorem]{Exercise}
\newtheorem{problem}[theorem]{Problem}
\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{note}[theorem]{Note}
\numberwithin{equation}{section}

\title{Hyers-Ulam stability of homomorphisms and $\ast$-homomorphisms on Hilbert $C^{*}$-modules}

\author[a]{Sajjad Khan}
\address[a]{Department of Mathematics, Hanyang University, Seoul 04763, Korea}
\ead{sajjadafridi@hanyang.ac.kr}

\author[b]{Choonkil Park\corref{mycorrespondingauthor}}
\address[b]{Research Institute for Convergence of Basic Sciences, Hanyang University,  Seoul 04763, Korea}
\ead{baak@hanyang.ac.kr}
\cortext[mycorrespondingauthor]{Corresponding author}

\begin{abstract}
In this paper, we introduce the idea of $\ast$-homomorphism on a Hilbert $C^{*}$-module. Furthermore, we prove the Hyers-Ulam stability of homomorphisms and $\ast$-homomorphisms on Hilbert $C^{*}$-modules using the fixed point method.
\end{abstract}

\begin{keyword} Hilbert $C^{*}$-module \sep Hilbert $C^{*}$-module homomorphism \sep Hilbert $C^{*}$-module $\ast$-homomorphism \sep fixed point method \sep Hyers-Ulam stability. 
\MSC[2020] 39B52 \sep 39B82  \sep 46L08  \sep 47B48  \sep 47H10
\end{keyword}
\end{frontmatter}

\section{Introduction and preliminaries} 

The stability problem of functional equations was arose from a question posed by Ulam \cite{Ulam} 
in 1940.
Specifically, the question was: Let $(G,\ast)$ be a group and $(G^{\diamond},\diamond, d^{\diamond})$ be a metric group with the metric $d^{\diamond}$. For any $\epsilon >0$, does there exist a $\delta >0$ such that if $f:G\rightarrow G^{\diamond}$ satisfies $d(f(x \ast y),f(x)\diamond f(y))< \delta$ for all $x,y \in G$, then a homomorphism $h:G \rightarrow G^{\diamond}$ exists with $d^{\diamond}(f(x),h(x))<\epsilon$ for all $x \in G$? If such $h$ exists, then the functional equation $h(x \ast y)=h(x)\diamond h(y)$ is stable. In 1941, Hyers \cite{Hyers} provided the first affirmative answer for approximately additive mappings in the context of Banach spaces. In 1978, Rassias \cite{Rassias 1978} generalized the Hyers' result for linear mappings. G\v{a}vruta \cite{Gavruta} introduced a general control function instead of the Cauchy difference introduced in \cite{Rassias 1978}, and further extended the Rassias' results. Rassias \cite{Rassias 1990} posed the question in 1990 about whether a similar theorem could be established for 
$p\geq 1$. Gajda \cite{Gajda} provided an affirmative answer to this question for 
$p > 1$, following the same approach as in \cite{Rassias 1978} by Rassias. Gajda \cite{Gajda}, along with Rassias and \v{S}emrl \cite{Rassias 1993}, showed that the Rassias' results cannot be established for $p=1$.\

In 1991, Baker \cite{Baker} used the fixed point method for the first time to study the stability of functional equations. Then, Radu \cite{Radu} and Cadariu and Radu \cite{Cadariu1} studied the stability of functional equations via fixed point method. In 1996, Isac and Rassias \cite{Isac} were the first to apply the stability theory of functional equations to establish new fixed point theorems and their applications. Using fixed point methods, many authors have extensively investigated the stability problem of various functional equations (see, for example \cite{Cadariu2, Siriluk, Park 2017}).

Park \cite{Park rho 1, Park rho 2} introduced additive $\rho$-functional inequalities and established the Hyers-Ulam stability of these inequalities in both Banach spaces and non-Archimedean Banach spaces. The stability problems of various functional equations, functional inequalities, and differential equations have been thoroughly studied by numerous researchers (see \cite{Sajjad1, Sajjad2, Park bi-derivation}).

In mathematics, the notion of a Hilbert $C^*$-module generalizes the concept of a Hilbert space by replacing the field of complex numbers with a general $C^{*}$-algebra. This concept was first introduced by Kaplansky in  \cite{kaplansky}. The stability of various mappings on Hilbert $C^{*}$-modules has been investigated by several mathematicians (see \cite{Ali, Ismail}).

Let $\mathcal{A}$ be a $C^{*}$-algebra and $\Xi$ be a complex linear space equipped with a compatible left $\mathcal{A}$-module action (i.e., $\lambda(a x)=(\lambda a)x=a(\lambda x)$ for $x \in \Xi$, $a \in \mathcal{A}$, $\lambda\in\mathbb{C}$). The space $\Xi$ is called a (left) pre-Hilbert $\mathcal{A}$-module or (left) inner product $\mathcal{A}$-module if there exists a mapping $\langle\cdot, \cdot\rangle: \Xi \times \Xi \rightarrow \mathcal{A}$, satisfying the following conditions: For all $x,y,z \in \Xi, \alpha,\beta \in \mathbb{C}, a \in \mathcal{A},$
\begin{itemize}
\item[(i)] $\langle \alpha x + \beta y,z  \rangle= \alpha \langle x,z \rangle + \beta \langle y,z \rangle$;
\item[(ii)] $\langle a x,y \rangle=a \langle x,y \rangle $;
\item[(iii)] $\langle x,y \rangle^{*} =\langle y,x \rangle$;
\item[(iv)] $\langle x,x \rangle\geq 0$ and $\langle x,x\rangle=0$ if and only if $x=0.$
\end{itemize}
The (left) pre-Hilbert space $\Xi$ is considered a (left) Hilbert $\mathcal{A}$-module or a (left) Hilbert $C^{*}$-module over the $C^{*}$-algebra $\mathcal{A}$ if it is complete with the induced norm $\|x\|:=\|\langle x , x\rangle\|^\frac{1}{2}$ \cite{Lance}. Right Hilbert $C^{*}$-modules can be defined in a similar way. In this paper, we will simply refer to a (left) Hilbert $C^*$-module over the $C^{*}$-algebra $\mathcal{A}$ as a Hilbert $C^*$-module. As an example, it is easy to see that every complex Hilbert space $\Xi$ is a Hilbert $C^*$-module over the $C^{*}$-algebra $\mathbb{C}$, with its inner product. Similarly, every $C^*$-algebra $\mathcal{A}$ can be regarded as a Hilbert $C^*$-module where, 
$\langle x,y\rangle = x y^{*}$, for $x,y \in \mathcal{A}$.

Khan \textit{et al.} \cite{Sajjad3} introduced the idea of fuzzy Hilbert $C^{*}$-module homomorphisms and fuzzy Hilbert $C^{*}$-module derivations and studied their Hyers-Ulam stability while in \cite{Sajjad4} they introduced Jordan homomorphisms and Jordan $\ast$-homomorphisms in Hilbert $C^{*}$-modules and discussed their Hyers-Ulam stability.

 In this paper, we introduce Hilbert $C^{*}$-module $\ast$-homomorphisms and prove the Hyers-Ulam stability of the Hilbert $C^{*}$-module homomorphisms and Hilbert $C^{*}$-module $\ast$-homomorphisms. It is observed that the definition of a Hilbert $C^{*}$-module homomorphism first appeared in \cite{Homomorphism definition}. 

\begin{defn} \cite{Homomorphism definition}
Let $\mathcal{A}$ and $\mathcal{B}$ be $C^{*}$-algebras, $\Xi$ be a Hilbert $\mathcal{A}$-module and $\nabla$ be a Hilbert $\mathcal{B}$-module. A mapping $H:\Xi \rightarrow \nabla$ is called a Hilbert $C^{*}$-module homomorphism if $H$ is $\mathbb{C}$-linear and satisfies
\begin{equation}\label{1.1}
H\left( \langle x,y\rangle z\right) =\langle H(x),H(y)\rangle H(z)
\end{equation}
for all $x,y,z \in \Xi $.
\end{defn}

\begin{defn}\label{1.2}
A Hilbert $C^{*}$-module homomorphism $H: \Xi \rightarrow \nabla $ is called a Hilbert $C^{*}$-module $\ast$-homomorphism if 
\begin{equation}
H\left( \langle x,y\rangle ^{*} z\right) =\langle H(x),H(y)\rangle ^{*} H(z)
\end{equation}
for all $x,y,z \in \Xi $.
\end{defn}

We recall a fundamental result in the fixed point theory. 
\begin{thm}\label{FPA}\cite{Cadariu1}
 Let (X,d) be a complete generalized metric space
and $J: X\rightarrow X$ be a strictly contractive mapping with
Lipschitz constant $L<1$. Then, for all $x\in X$, either
\begin{equation*}
d(J^n x,J^{n+1}x)=\infty
\end{equation*}
for all nonnegative integers $n$ or there exists a positive
integer $n_0$ such that

$(1)$ $d(J^n x,J^{n+1}x)<\infty$ for all $n\geq n_0$;

$(2)$ the sequence $\{J^n x\}$ converges to a fixed point $y^*$ of
$J$;

$(3)$ $y^*$ is the unique fixed point of $J$ in the set $Y=\{y\in
X : d(J^{n_0}x,y)<\infty\}$;

$(4)$ $d(y,y^*)\leq \frac{1}{1-L}d(y,Jy)$ for all $y\in Y$.
\end{thm}

Throughout the paper, we will denote $\Xi$ as a Hilbert $\mathcal{A}$-module and $\nabla$ as a Hilbert $\mathcal{B}$-module. The paper is organized as follows: In Section 2, we prove the Hyers-Ulam stability of Hilbert $C^{*}$-module homomorphisms. In Section 3, we establish the same results for Hilbert $C^{*}$-module $\ast$-homomorphisms. 
%




\section{Stability of Hilbert $C^{*}$-module homomorphisms}
We will need the following lemma for our main results.

\begin{lem}\label{c-linear}\cite{c-linear}
Let $X$ and $Y$ be linear spaces and $f:X \rightarrow Y$ be an additive mapping such that $f(\mu x)=\mu f(x)$ for all $x \in X$ and any $\mu \in \mathbb{T}=\lbrace z \in \mathbb{C} : \vert z\vert =1\rbrace$. Then, the mapping f is $\mathbb{C}$-linear.
\end{lem}

\begin{thm}\label{theorem 1}
Let $f:\Xi \rightarrow \nabla$ be a mapping for which there exists a function  $\varphi: \Xi^{3}\rightarrow [0,\infty)$ such that 
\begin{equation}\label{2.1}
\vert\vert f(\mu x +y)-\mu f(x)-f(y)\vert\vert \leq \varphi(x,y,0),
\end{equation}
\begin{equation}\label{2.2}
\| f(\langle x,y\rangle z)-\langle f(x), f(y) \rangle f(z) \| \leq \varphi(x,y,z)
\end{equation}
for all $x,y,z \in \Xi$ and all $\mu \in \mathbb{T}$. If there exists  $0\leq L < 1$ such that $\varphi (x,y,z)\leq 2 L \varphi (\frac{x}{2},\frac{y}{2},\frac{z}{2})$ for all $x,y,z \in \Xi $, then there exists a unique Hilbert $C^{*}$-module homomorphism $H:\Xi \rightarrow \nabla$ such that 
\begin{equation}\label{2.3}
\| f(x)- H(x) \| \leq \dfrac{1}{2-2L}\varphi(x,x,0)
\end{equation}
for all $x \in \Xi$.
\end{thm}

\begin{proof}
Consider $\mathcal{S}=\lbrace g:\Xi\rightarrow \nabla \rbrace$, and define the generalized metric on $\mathcal{S}$:
\begin{equation}\label{2.4}
d(g,h):= \text{inf} \lbrace c \in (0,\infty):\|g(x)-h(x)\|\leq c \varphi(x,x,0), \forall x \in \Xi\rbrace.
\end{equation}
It is easy to show that $(\mathcal{S},d)$ is a complete metric space (see \cite{m-r}). 

Now, consider the linear mapping $J: \mathcal{S}\rightarrow \mathcal{S}$
\begin{equation}\nonumber
(Jg)(x):= \dfrac{1}{2}g(2x)
\end{equation}
for all $x\in \Xi$.\\
By \cite[Theorem 3.1]{Cadariu1}, $J$ is a strictly contractive mapping.

Letting $x=y$, $\mu =1$ in (\ref{2.1}), we obtain
\begin{equation}\label{2.5}
\left \| f(2x)-2f(x) \right \|\leq \varphi(x,x,0)
\end{equation}
for all $x \in \Xi$. This implies that 
\begin{equation}\nonumber
\left \| f(x)- \dfrac{1}{2}f(2x) \right \| \leq\frac{1}{2}\varphi(x,x,0)
\end{equation}
for all $x \in \Xi$. Thus $d(f,Jf)\leq \frac{1}{2}$.

By Theorem \ref{FPA}, there exists a mapping $H:\Xi \rightarrow \nabla$ which is a fixed point of $J$ such that 
\begin{equation}\label{2.6}
H(2x)=2H(x)
\end{equation}
for all $x \in \Xi$ and $\displaystyle{\lim_{n\to\infty}} d(J^{n}f,H)=0$. This implies that 
\begin{equation}\label{2.7}
 \lim_{n\to\infty}\dfrac{f(2^{n}x)}{2^{n}}=H(x) 
\end{equation}
for all $x \in \Xi.$ Notice that  $H$ is the unique fixed point of $J$ in the set 
\begin{equation}\nonumber
U=\lbrace g \in \mathcal{S}: d(g,f)<\infty\rbrace.
\end{equation}
This implies that $H$ is the unique fixed point satisfying (\ref{2.6}) such that there exists $c \in (0,\infty)$ satisfying 
\begin{equation}\nonumber
\| f(x) - H(x)\|\leq c \varphi(x,x,0)
\end{equation}
for all $x \in \Xi$.\\
Applying Theorem \ref{FPA} once again, we obtain $d(f,H)\leq \frac{1}{1-L}d(f,Jf)\leq \frac{1}{2-2L}.$ This implies 
\begin{equation}\nonumber
\|f(x)-H(x)\|\leq \frac{1}{2-2L} \varphi(x,x,0)
\end{equation}
for all $x\in \Xi$ and (\ref{2.3}) holds.

Now we show that $H$ is $\mathbb{C}$-linear. From the assumption $\varphi(x,y,z) \leq 2L \varphi(\frac{x}{2},\frac{y}{2},\frac{z}{2})$ for all $x,y,z \in \Xi$, it follows that
\begin{equation}\label{2.8}
\lim_{n\to\infty} \frac{1}{2^{n}}\varphi (2^{n}x,2^{n}y,2^{n}z)\leq \lim_{n\to\infty} \frac{2^{n}L^{n}}{2^{n}}\varphi (x,y,z)=0.
\end{equation}
In (\ref{2.1}), replacing $x$ by $2^{n}x$ and $y$ by $2^{n}y$, respectively, we have
\begin{equation}\nonumber
\|f(\mu 2^{n}x+2^{n}y)-\mu f(2^{n}x)-f(2^{n}y)\|\leq \varphi(2^{n}x,2^{n}y,0).
\end{equation}
Taking the limit as $n \rightarrow \infty$, using (\ref{2.7}) and (\ref{2.8}), we obtain 
\begin{equation}\label{2.9}
H(\mu x+y)=\mu H(x)+H(y)
\end{equation}
for all $\mu \in \mathbb{T}$ and all $x,y \in \Xi.$ The case  $\mu=1$ in (\ref{2.9}) implies that $H$ is additive and so $H(0)=0$, together with these, Lemma \ref{c-linear} implies that $H$ is $\mathbb{C}$-linear.

Putting $x,y,z$ as $2^{n}x,2^{n}y,2^{n}z$, respectively, in (\ref{2.2}), we obtain 
\begin{equation}\nonumber 
\|f(\left\langle  2^{n}x,2^{n}y \right\rangle  2^{n}z)- \left\langle  f(2^{n}x),f(2^{n}y)\right\rangle  f(2^{n}z)\|\leq \varphi(2^{n}x,2^{n}y,2^{n}z).
\end{equation}
This further implies that 
\begin{align*}\nonumber
\left \| 2^{-3n}f(2^{3n}\left\langle  x,y\right\rangle  z)-\left\langle  \frac{f(2^{n}x)}{2^{n}},\frac{f(2^{n}y)}{2^{n}}\right\rangle  \frac{f(2^{n}z)}{2^{n}} \right \| & \leq 2^{-3n}\varphi(2^{n}x,2^{n}y,2^{n}z)\\
& \leq  2^{-3n}8^{n}L^{n}\varphi(x,y,z)
\end{align*}
for all $x,y,z \in \Xi$, since $\varphi(x,y,z)\leq 2L \varphi(\frac{x}{2},\frac{y}{2},\frac{z}{2}) \leq 8L \varphi(\frac{x}{2},\frac{y}{2},\frac{z}{2})$.\\
Taking the limit as $n\rightarrow \infty $, applying (\ref{2.7}) and (\ref{2.8}), we get 
\begin{equation}\nonumber
H(\left\langle  x,y\right\rangle  z) = \left\langle H(x), H(y)\right\rangle H(z)
\end{equation}
for all $x,y,z \in \Xi$. Thus, $H$ is a Hilbert $C^{*}$-module homomorphism. 
\end{proof}

The following corollary gives us the Hyers-Ulam-Rassias stability of homomorphisms in $\Xi$.

\begin{cor}
Let $p\in [0,1)$, $\theta$ be a nonnegative real number and $f:\Xi \rightarrow \nabla $ be a mapping such that 
\begin{equation}\label{2.10}
\| f(\mu x +y)-\mu f(x)-f(y)\|  \leq \theta \left( \|x\|^{p}+\|y\|^{p}\right), 
\end{equation}
\begin{equation}\label{2.11}
\| f(\left\langle  x,y\right\rangle z)- \left\langle f(x),f(y)\right\rangle f(z) \|  \leq \theta \left( \|x\|^{p}+\|y\|^{p}+\|z\|^{p} \right) 
\end{equation}
for all $x,y,z \in \Xi$ and all $\mu \in \mathbb{T}$. Then there exists a unique Hilbert $C^{*}$-module homomorphism $H:\Xi \rightarrow \nabla $ such that 
\begin{equation}
\| f(x)- H(x) \| \leq \frac{2 \theta}{2-2^{p}}\|x\|^{p}
\end{equation}
for all $x \in \Xi$.
\end{cor}

\begin{proof}
Take $\varphi(x,y,z):=\theta\left( \|x\|^{p}+\|y\|^{p}+\|z\|^{p}\right)$ and $L=2^{p-1}$. Then the proof follows from Theorem \ref{theorem 1}.
\end{proof}

\begin{thm}\label{theorem 2}
Let $f:\Xi \rightarrow\nabla $ be a mapping for which there exists a function $\varphi: \Xi^{3}\rightarrow [0,\infty)$ satisfying (\ref{2.1}) and (\ref{2.2}) for all $x,y,z \in \Xi$ and all $\mu \in \mathbb{T}$. 
If there exists  $0 \leq L < 1$ such that $\varphi (x,y,z)\leq \frac{L}{8} \varphi (2x,2y,2z)$ for all $x,y,z \in \Xi$, then there exists a unique Hilbert $C^{*}$-module homomorphism $H:\Xi\rightarrow \nabla$ such that 
\begin{equation}\label{2.13}
\| f(x)- H(x) \| \leq \dfrac{L}{8-8L}\varphi(x,x,0)
\end{equation}
for all $x \in \Xi$.
\end{thm}

\begin{proof}
Consider the linear mapping $J: \mathcal{S}\rightarrow \mathcal{S}$
\begin{equation}\nonumber
Jg\left( x \right) := 2g\left( \frac{1}{2}x\right) 
\end{equation}
for all $x\in \Xi$.\\
It follows from (\ref{2.5}) that 
\begin{equation}\nonumber
\left \| f(2x)-2f\left( \frac{x}{2}\right)  \right \| \leq \varphi\left( \frac{x}{2},\frac{x}{2},0\right)  \leq \frac{L}{8} \varphi(x,x,0)
\end{equation}
for all $x \in \Xi$. By (\ref{2.4}), we have $d(f,Jf)\leq \frac{L}{8}$.

By Theorem \ref{FPA}, there exists a mapping $H:\Xi \rightarrow \nabla$ which is a fixed point of $J$ such that 
\begin{equation}\label{2.14}
H\left( \frac{1}{2}x\right) =\frac{1}{2}H(x)
\end{equation}
for all $x \in \Xi$. Also $\displaystyle{\lim_{n\to\infty}}d(J^{n}f,H)=0$. This implies that 
\begin{equation}\label{2.15}
 \lim_{n\to\infty}2^{n}f\left( \frac{x}{2^{n}}\right) =H\left( x\right) 
\end{equation}
for all $x \in \Xi.$ Notice that  $H$ is the unique fixed point of $J$ in the set 
\begin{equation}\nonumber
U=\lbrace g \in \mathcal{S}: d(g,f)<\infty\rbrace.
\end{equation}
This implies that $H$ is the unique fixed point satisfying (\ref{2.14}) such that there exists $c \in (0,\infty)$ satisfying 
\begin{equation}
\| f(x) - H(x)\|\leq c \varphi(x,x,0)
\end{equation}
for all $x \in \Xi$.\\
Applying Theorem \ref{FPA} once again, we obtain $d(f,H)\leq \frac{1}{1-L}d(f,Jf)\leq \frac{L}{8-8L}.$ This implies 
\begin{equation}\nonumber
 \|f(x)-H(x) \|\leq \frac{L}{8-8L} \varphi(x,x,0)
\end{equation}
for all $x\in \Xi$ and (\ref{2.13}) holds.

From the assumption $\varphi(x,y,z) \leq \frac{L}{8}\varphi(2x,2y,2z)\leq\frac{L}{2}\varphi(2x,2y,2z) $ for all $x,y,z \in \Xi$, it follows that 
\begin{equation}\label{2.17}
\lim_{n\to\infty} 2^{n}\varphi \left( \frac{x}{2^{n}},\frac{y}{2^{n}},\frac{z}{2^{n}}\right) \leq \lim_{n\to\infty} \frac{2^{n}L^{n}}{2^{n}}\varphi (x,y,z)=0.
\end{equation}
In (\ref{2.1}), replacing $x$ by $\frac{x}{2^{n}}$, $y$ by $\frac{y}{2^{n}}$, multiplying both side by $2^{n}$, letting $n \rightarrow \infty$, using (\ref{2.15}) and (\ref{2.17}), we obtain
\begin{equation}\label{2.18}
H(\mu x+y)=\mu H(x)+H(y)
\end{equation}
for all $x,y \in \Xi$ and all $\mu \in \mathbb{T}$. The case $\mu =1$ in (\ref{2.18}) implies that $H$ is additive and by Lemma \ref{c-linear}, $H$ is $\mathbb{C}$-linear.

Replacing $x,y,z$ with $\frac{x}{2^{n}},\frac{y}{2^{n}},\frac{z}{2^{n}}$, respectively, in (\ref{2.2}), we obtain 
\begin{align*}\nonumber
\left \| 2^{3n} f \left(\left\langle  \frac{x}{2^{n}},\frac{y}{2^{n}} \right\rangle \frac{z}{2^{n}} \right) - \left\langle 2^{n}f \left( \frac{x}{2^{n}} \right) ,2^{n} f \left( \frac{y}{2^{n}} \right) \right\rangle 2^{n} f \left( \frac{z}{2^{n}}\right) \right \|  & \leq 2^{3n} \varphi \left( \frac{x}{2^{n}},\frac{y}{2^{n}},\frac{z}{2^{n}}\right)\\
& \leq 2^{3n}2^{-3n}L^{n} \varphi \left( x,y,z \right)
\end{align*}
for all $x,y,x \in \Xi$. Taking $n \rightarrow \infty$, using (\ref{2.15}) and (\ref{2.17}), we obtain 
\begin{equation}\nonumber
H(\left\langle x,y\right\rangle z)=\left\langle H(x),H(y)\right\rangle H(z)
\end{equation}
for all $x,y,z \in \Xi$. Thus $H$ is a Hilbert $C^{*}$-module homomorphism. 
\end{proof}

\begin{cor}
Let $p>3$, $\theta$ be a nonnegative real number and $f:\Xi \rightarrow \nabla$ be a mapping satisfying (\ref{2.10}) and (\ref{2.11}). Then there exists a unique Hilbert $C^{*}$-module homomorphism $H: \Xi \rightarrow \nabla$ such that 
\begin{equation}\nonumber
\| f(x)- H(x) \| \leq \dfrac{2\theta}{2^{p}-2^{3}}\|x\|^{p}
\end{equation}
for all $x \in \Xi$.
\end{cor}

\begin{proof}
Take $\varphi(x,y,z):=\theta\left( \|x\|^{p}+\|y\|^{p}+\|z\|^{p}\right)$ and $L=2^{3-p}$. Then the proof follows from Theorem \ref{theorem 2}.
\end{proof}

\section{Stability of Hilbert $C^{*}$-module $\ast$-homomorphisms} 

In this section, we establish the Hyers-Ulam stability of Hilbert $C^{*}$-module $\ast$-homomorphisms. 

\begin{theorem}\label{theorem 3}
Let $f:\Xi \rightarrow \nabla$ be a mapping for which there exists a function $\varphi: \Xi^{3}\rightarrow [0,\infty)$ satisfying (\ref{2.1}) and (\ref{2.2}) and 
\begin{equation}\label{3.1}
\| f(\left\langle x,y\right\rangle ^{*}z)- \left\langle f(x),f(y)\right\rangle ^{*}f(z) \| \leq \varphi (x,y,z) 
\end{equation}
for all $x,y,z \in \Xi$. If there exists  $0\leq L < 1$ such that $\varphi (x,y,z)\leq 2 L \varphi (\frac{x}{2},\frac{y}{2},\frac{z}{2})$ for all $x,y,z \in \Xi$, then there exists a unique Hilbert $C^{*}$-module $*$-homomorphism $H:\Xi \rightarrow \nabla $ such that 
\begin{equation}\label{3.2}
\| f(x)- H(x) \| \leq \dfrac{1}{2-2L}\varphi(x,x,0)
\end{equation}
for all $x \in \Xi$.
\end{theorem}
\begin{proof}
From Theorem \ref{theorem 1}, there exists a unique Hilbert $C^{*}$-module homomorphism $H$ satisfying (\ref{3.2}) and defined as 
\begin{equation}\label{3.3}
\lim_{n\to\infty}\frac{f(2^{n}x)}{2^{n}} =H\left( x \right) 
\end{equation}
for all $x \in \Xi$.

Replacing $x,y,z$ by $2^{n}x,2^{n}y,2^{z}$, respectively, in (\ref{3.1}), we obtain 
\begin{equation}\nonumber
\| f (\left\langle 2^{n}x,2^{n}y\right\rangle^{*} 2^{n}z) - \left\langle f(2^{n}x),f(2^{n}y)\right\rangle^{*} f(2^{n}z)\| \leq \varphi (2^{n}x,2^{n}y,2^{n}z).
\end{equation}
This further implies that 
\begin{align*}\nonumber 
\left \|2^{-3n} f (2^{3n}\left\langle x,y\right\rangle^{*} z) - \left\langle \frac{f(2^{n}x)}{2^{n}},\frac{f(2^{n}y)}{2^{n}}\right\rangle^{*} \frac{f(2^{n}z)}{2^{n}} \right \| & \leq 2^{-3n} \varphi (2^{n}x,2^{n}y,2^{n}z)\\
& \leq 2^{-3n}2^{3n}L^{n} \varphi (x,y,z)
\end{align*}
for all $x,y,z\in \Xi$. Taking the limit as $n \rightarrow \infty$, using (\ref{2.15}) and (\ref{2.17}), we obtain
\begin{equation}\nonumber 
H(\left\langle x,y\right\rangle ^{*}z)=\left\langle H(x),H(y)\right\rangle^{*} H(z)
\end{equation}
for all $x,y,z \in \Xi$. So $H$ is a Hilbert $C^{*}$-module $\ast$-homomorphism.
\end{proof}

The following corollary gives us the Hyers-Ulam-Rassias stability of Hilbert $C^{*}$-module $\ast$-homomorphisms.

\begin{cor}
Let $p\in [0,1)$, $\theta$ be a nonnegative real number and $f:\Xi \rightarrow \nabla$ be a mapping satisfying (\ref{2.10}), (\ref{2.11}) and 
\begin{equation}\label{3.4}
\| f(\left\langle x,y\right\rangle ^{*}z)- \left\langle f(x),f(y)\right\rangle ^{*}f(z) \| \leq \theta (\|x\|^{p}+\|y\|^{p}+\|z\|^{p})
\end{equation}
for all $x,y,z \in \Xi$. Then there exists a unique Hilbert $C^{*}$-module $\ast$-homomorphism $H:\Xi \rightarrow \nabla$ such that 
\begin{equation}\nonumber
\| f(x)- H(x) \| \leq \frac{2 \theta}{2-2^{p}}\|x\|^{p}
\end{equation}
for all $x \in \Xi$.
\end{cor}

\begin{proof}
The proof follows from Theorem \ref{theorem 3} by taking $\varphi (x,y,z)=\theta (\|x\|^{p}+\|y\|^{p}+\|z\|^{p})$ and $L=2^{p-1}$. 
\end{proof}

\begin{theorem}\label{theorem 4}
Let $f:\Xi \rightarrow \nabla$ be a mapping for which there exists a function $\varphi: \Xi^{3}\rightarrow [0,\infty)$ satisfying (\ref{2.1}) and (\ref{2.2}) and (\ref{3.1}). If there exists  $0\leq L < 1$ such that $\varphi (x,y,z)\leq \frac{L}{8} \varphi (2x,2y,2z)$ for all $x,y,z \in \Xi$, then there exists a unique Hilbert $C^{*}$-module $\ast$-homomorphism $H:\Xi \rightarrow \nabla$ such that 
\begin{equation}\label{3.5}
\| f(x)- H(x) \| \leq \dfrac{L}{8-8L}\varphi(x,x,0)
\end{equation}
for all $x \in \Xi$.
\end{theorem}
\begin{proof}
By Theorems \ref{theorem 2}, there exists a unique Hilbert $C^{*}$-module homomorphism $H:\Xi\rightarrow \nabla$ satisfying (\ref{3.5}) and defined as
\begin{equation}\label{3.6}
\lim_{n\to\infty}2^{n}f\left( \frac{x}{2^{n}}\right)  =H \left( x \right) 
\end{equation}
for all $x \in \Xi$.\\
Replacing $x,y,z$ by $\frac{x}{2^{n}},\frac{y}{2^{n}},\frac{z}{2^{n}}$, respectively, in (\ref{3.1}), we obtain
\begin{equation}\nonumber
\left \| 2^{3n}f\left( \left\langle \frac{x}{2^{n}},\frac{y}{2^{n}}\right\rangle ^{*}\frac{z}{2^{n}}\right) -\left\langle 2^{n}f\left( \frac{x}{2^{n}}\right) ,2^{n}f\left( \frac{y}{2^{n}}\right)  \right\rangle ^{*} 2^{n}f\left( \frac{z}{2^{n}} \right) \right \| \leq 2^{3n} \varphi \left( \frac{x}{2^{n}},\frac{y}{2^{n}},\frac{z}{2^{n}}\right) 
\end{equation}
for all $x,y,z \in \Xi$. Taking the limit as $n \rightarrow \infty$. using (\ref{2.15}) and (\ref{2.17}), we obtain
\begin{equation}\nonumber
H(\left\langle x,y\right\rangle ^{*}z)=\left\langle H(x),H(y)\right\rangle ^{*}H(z)
\end{equation}
for all $x,y,z \in \Xi$. So $H$ is a Hilbert $C^{*}$-module $\ast$-homomorphism.
\end{proof}

The following corollary gives us the Hyers-Ulam-Rassias stability of $\ast$-homomorphisms in $\Xi$.

\begin{cor}
Let $p>3$, $\theta$ be a nonnegative real number and $f:\Xi \rightarrow \nabla$ be a mapping  satisfying (\ref{2.10}), (\ref{2.11}) and (\ref{3.4}). Then there exists a unique Hilbert $C^{*}$-module $\ast$-homomorphism $H:\Xi\rightarrow \nabla$ such that 
\begin{equation}
\|f(x)-H(x)\| \leq \frac{2\theta}{2^{p}-2^{3}}\|x\|^{p} 
\end{equation}
for all $x \in \Xi$.
\end{cor}

\begin{proof}
Take $\varphi (x,y,z)= \theta (\|x\|^{p}+\|y\|^{p}+\|z\|^{p})$ and $L=2^{3-p}$. Then the proof follows from Theorem \ref{theorem 4}.
\end{proof}

\section{Conclusion}

In this paper, we introduced the idea of $\ast$-homomorphisms on Hilbert $C^{*}$-modules. Furthermore, we established the Hyers-Ulam and Hyers-Ulam-Rassias stability of Hilbert $C^{*}$-module homomorphisms and Hilbert $C^{*}$-module $\ast$-homomorphisms using the fixed point method.

\medskip

\section*{Declarations}

\medskip

\noindent \textbf{Availability of data and materials}\newline
\noindent No datasets were generated or analyzed during the current study.

\medskip

\noindent \textbf{Human and animal rights}\newline
\noindent We would like to mention that this article does not contain any studies
with animals and does not involve any studies over human being.

\medskip

\noindent \textbf{Conflict of interest}\newline
\noindent The authors declare that they have no competing interests.

\medskip

\noindent \textbf{Fundings} \newline
\noindent  Not applicable.

\medskip

\noindent \textbf{Authors' contributions}\newline
\noindent The authors equally conceived of the study, participated in its
design and coordination, drafted the manuscript, participated in the
sequence alignment, and read and approved the final manuscript. 

\medskip

{\footnotesize

\end{document}